# Adaptive Reinforcement Learning Planning: Harnessing Large Language Models for Complex Information Extraction

### v1

@@ 20             Replace target-net parameters $\theta^{\prime}$ = $\theta$ every $k$ steps
Furthermore, for different datasets of the same language and task type (such as different English RE datasets), we only need to mix the training data to train one decision model, which simplifies the training process of the decision model and improves the generalization of our method.
3.5 LLM for Relation/Event Classification
In addition, in some datasets, the relation/event types of the originally labeled data are not natural language expressions. Consequently, it is challenging for LLMs to comprehend the semantics of these relation/event words without undergoing fine-tuning training. To address this, we undertake a natural language conversion for these labels, such as converting "people-deceased_person-place_of_death" in NYT10 dataset to "dead in". Please refer to Appendix B for details, and Appendix C for prompt template.
3.6 LLM-Based Argument Extractor
We adopt an LLM-based extractor to solve the argument extraction task with different extraction orders. For EE tasks, the roles are determined by the event type, and the event type has a corresponding roles schema. The extractor extracts the arguments for the selected role. For RE tasks, the intuitive approach is to set two roles for all relation types: subject and object. However, unlike roles in EE, the words "subject" and "object" do not contain any information about the entity type, which sometimes leads to ambiguity. For example, the subject-object of the relation "affiliated" is person-organization in some datasets, while in other datasets it is department-organization. Therefore, we design the roles in RE tasks as "subject/object: entity type". The entity type is determined by the semantics of the selected relation type in the specific dataset. See Appendix B for examples.
The input of the arguments extractor is a dictionary, including the original sentence, the current relation/event type, the extracted content (role-argument pairs), and the role (argument type) to extract. At each step of argument extraction, the extractor returns the arguments corresponding to the specified role in a string, separated by commas. The prompt template is provided in Appendix C.
4 Experiments
4.1 Datasets
We evaluate our methods on several public and accessible complicated information extraction datasets, including the relation extraction datasets NYT, NYT11-HRL, Wiki80, SKE21, HacRED and DuIE, and the event extraction dataset DuEE and ACE05. These datasets are challenging for LLMs and widely used for evaluation of existing extraction methods. A brief introduction to these datasets is provided in Appendix A.
4.2 Comparing Methods and Metrics
We select some recently popular and high-performance LLMs as extractors in our experiments. We choose the open-source model mistral7B-instruct-v0.3333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Qwen1.5 Bai et al. (2023), and the famous GPT3.5-turbo Brown et al. (2020); Ouyang et al. (2022) for both English and Chinese datasets. We apply two previous multi-stage extraction methods to our testset, ChatIE Wei et al. (2023) and LLM(without fine-tuning)+Filter Ding et al. (2024) and compare our method with them. We also compare our method with random/fixed selection of actions, and specifically calculate the metrics in cases involving multiple triples or roles.
We design a relaxed evaluation method based on the exact match. We consider the correctness of the arguments at the word level and calculate the similarity between the extraction result and ground truth for each argument (according to Eq. 3). In the level of the entire triple or event, we believe the extraction result matches the ground truth only if the relation/event type is classified correctly and the word-level similarity is greater than the threshold for all arguments. We calculate and report precision (Prec.), recall (Reca.), and F1 scores for all the experiments according to the method above. We repeat the experiment 3 times and averaged the metrics as final results. See Appendix D for details.
4.3 Effectiveness of LLM Classification
We first assess the efficacy of LLM-based relation/event classification on the SKE21, DuEE and ACE05 datasets. Table 4 in Appendix E shows that LLMs classify the relation/event accurately and achieve stable and accurate results, setting the stage for the argument extraction.
@@ 4.3 Effectiveness of LLM Classification
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo – ChatIE 81.61 48.51 60.85 71.67 43.31 53.99 70.13 66.13 68.07 57.05 78.13 65.95
– Filter 22.34 25.89 23.98 43.08 51.27 46.82 58.38 72.56 64.70 - - -
@@
Qwen1.5-14B – ChatIE 55.99 19.85 29.31 48.94 35.25 40.99 47.40 39.14 42.87 49.19 38.06 42.91
– Filter 22.34 25.89 23.98 39.86 57.11 46.95 26.42 48.08 34.10 - - -
@@
Mistral-7B-v0.3 – ChatIE 34.56 7.09 11.76 31.50 8.73 13.67 42.84 31.18 36.09 37.73 11.94 18.14
– Filter 6.03 14.37 8.49 28.81 38.60 32.99 36.53 37.27 36.90 - - -
@@
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo – ChatIE 46.27 42.70 44.41 32.12 42.97 36.76 32.14 29.57 30.80 57.52 22.18 32.02
– Filter 39.12 44.06 41.44 33.06 47.19 38.88 17.08 18.15 17.60 - - -
@@
Qwen1.5-14B – ChatIE 39.66 15.66 22.45 27.48 19.46 22.78 16.67 18.29 17.44 41.38 12.29 18.95
– Filter 12.36 19.22 15.04 25.82 29.97 27.74 6.64 9.50 7.82 - - -
@@
Mistral-7B-v0.3 – ChatIE 25.61 11.84 16.19 19.38 13.51 15.92 6.18 9.14 7.37 51.16 47.51 49.27
– Filter 38.50 59.91 46.88 16.67 19.41 17.94 8.34 9.92 9.06 - - -
@@
4.4 Main Results
Table 1 and  2 show the extraction results on the Chinese and English datasets respectively. ChatIE Wei et al. (2023) uses a multi-step instruction to guide chatgpt to extract, we also apply this instruction template to the open source model and report the results, as shown in the first row of each block. Ding et al. (2024) adds the filtered results of small models to the prompt of LLMs, and we report the performance of this method on our test set, as shown in the second row of each block. All the LLMs in the table have not been fine-tuned on the labeled train set.
@@ 4.4 Main Results
4.5 Results of Different Step Orders
In addition, we also compared the extraction results of different order selection methods for entity extraction, using Qwen for Chinese data and Mistral for English data, as shown in Table 1 and Table 2. In most cases, RL-based extraction order can achieve the best results, and the metrics of fixed extraction order are lower than random order. These results imply that the optimal order varies across distinct instances, and our method could adaptively make the best plan.
4.6 Complicated Extraction Settings
@@
The results of Table 5 in Appendix E show that when the IE task becomes complicated, the extraction performance of ChatIE is far inferior to our method. In addition, compared with random or fixed extraction order, our method achieves significant improvements in the metrics. This also suggests that when LLMs undertake extraction tasks in complex scenarios, an appropriate order can further augment the model’s extraction ability.
5 Conclusion
@@
Limitations
Despite the effectiveness of our model, there are still some limitations of our method.
First, our reward model is concise and effective. In most cases, it is not difficult for a large model to make binary reward assignments based on the semantics of extraction results and the ground-truth. However, it only considers the correctness of each step and lacks the evaluation of the complete extraction results of the entire sentence. Subsequent research can further explore the design of rewards of LLM-based environments.
Second, the decision module of sequential argument extraction in our framework is still based on small models. Furthermore, the training of decision model requires extensive interactions with LLMs, which is relatively time-consuming. In future research, we could contemplate merging similar datasets or texts to train the decision model, or further explore approaches for LLMs to directly undertake the order-decision tasks.
Ethics Statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Human annotations are only used in methodological research at the beginning of the work, to assist in analyzing the feasibility of the proposed solution. Annotators consented to the use of data for research purposes. We ensure that the privacy of all annotators is protected throughout the annotation process, and all of them are adequately paid according to local standards. Human annotations are not applied during the evaluation of our method.
Risks
In this paper, all datasets are obtained from public sources. The datasets adopted have been anonymized and do not contain offensive information. However, we cannot guarantee that the datasets do not contain socially harmful or toxic language.
Appendix A Dataset Introduction
NYT series
NYT is based on the articles in New York Times. There are many derived datasets with better labeling. NYT10 Riedel et al. (2010) and NYT11 Hoffmann et al. (2011) label the complete entities. Moreover, NYT10-HRL and NYT11-HRL Takanobu et al. (2019) are better versions that are processed by optimizing the relation labels.
Wiki80
Wiki80 Han et al. (2019) is a sentence-level dataset based on the few-shot dataset FewRel Han et al. (2018), which contains 80 types of relations. The number of each relation is 700, with a total of 56,000 samples.
HacRED
HacRED Cheng et al. (2021) [^4] https://github.com/qiaojiim/hacred is a novel challenging extraction dataset. It analyzes the performance gap between popular datasets and practical applications, and carefully selects and designs more hard cases. HacRED consists of 65,225 relational facts annotated from 9,231 wiki documents with sufficient and diverse hard cases, which poses a very high challenge to many current complex extraction methods.
SKE21
SKE19555http://ai.baidu.com/broad/download?dataset=sked is published by Baidu, and is currently the largest dataset available for complex relational triple extraction. Since its testing set is unpublished, and there are some errors in the validation set, a version named SKE21 is published by Xie et al. (2021). The testing set of SKE21 is carefully manually relabeled and contains 1,150 sentences and 2,765 annotated triples.
DuIE
DuIE Li et al. (2019) is a Chinese information extraction dataset. It contains more than 210, 000 sentences and 48 pre-defined schema gathered from Baidu Encyclopedia, Baidu Tieba and Baidu Information Stream. Compared to the previous version, it contains 5 Multiple-O-values schema (Schema with multiple object slots and values), which greatly increase the difficulty of the task.
DuEE
DuEE Li et al. (2020) is a Chinese document-level event extraction dataset. It contains 11, 224 documents categorized into 65 event types, along with 41, 520 event arguments mapped to 121 argument roles, which is the largest Chinese EE dataset.
ACE05
ACE 2005666https://catalog.ldc.upenn.edu/LDC2006T06 Multilingual Training Corpus was developed by the Linguistic Data Consortium (LDC) and contains approximately 1,800 files of mixed genre text in English, Arabic, and Chinese annotated for entities, relations, and events. In this paper we use the English events corpus, it provides event annotations in document and sentence levels from a variety of domains such as newswires and online forums.
Appendix B The Inputs of LLMs
In Sections 3.5 and  3.6, we introduce the inputs for the LLM-based classifier and extractor respectively. We describe the natural language conversion of relation/event types, and the design of roles in RE tasks when extracting arguments. Here we provide some illustrative intuitive examples based on the NYT11 dataset.
B.1 Natural Language Conversion of Types
We convert the original relation words in the dataset into natural language expressions that are conducive to the comprehension of LLMs. For example,
    /location/location/contains -> location contains
    /business/person/company -> work at
    /people/person/place-of-birth -> born in
    /business/company/founders -> founder
B.2 Roles in RE Tasks
We design the roles in RE tasks as "subject/object: entity type" to reduce the ambiguity of relation types and promote the argument extraction of LLMs. For example, in NYT11 datasets:
    the roles of relation "location contains" are: "subject: place" and "object: place"
    the roles of relation "work at" are: "subject: person" and "object: company"
    the roles of relation "born in" are: "subject: person" and "object: place"
    the roles of relation "founder" are: "subject: company" and "object: person"
Appendix C Prompt Templates
Prompt template for classification:
Given the origin sentence, and a relation/event type list $L$, Please select the relation/event contained in this sentence from the list $L$ If more than one relation/event in $L$ is included in this sentence, please separate them with commas(’,’) and output.
Here are some examples:
Input: $s_{1},L_{1}$
Output: $o_{1}$
Now given the following sentence and relation/event list, please select all relations or events contained in the sentence, and output them.
Note: Consider only the types provided in the list.
Prompt template for entity extraction:
Given the text $s$ and the interested relation/event $p$, we may also give the extracted content $C$. Please identify the argument of the role $r$ according to the given content.
Here are some examples:
Input: {$s_{1},p_{1},C_{1}$, "the role I want": $r_{1}$}
Output: $a_{1}$
Now given the following input, please complete the extracting task.
Note: The output content should only contains the arguments of the role I want, and needs to match the relation/event type and extracted content provided in the input. If there are multiple corresponding arguments, please separate them with commas(’,’) and output.
Prompt template for reward assignment:
Now there is a model extraction competition. Please judge whether the answer extracted by the model is correct based on the ground-truth and give a score. If the extraction is correct, please give a score of 1, otherwise 0.
Here are some examples:
Extracted: USA; Ground-truth: the United States; Output: 1
Extracted: Beijing City; Ground-truth: Nanjing City; Output: 0
You only need to output the score value 1 or 0, without any redundant output.
Note: The extraction result does not have to be exactly the same as the ground-truth. It can be judged as correct if the meaning is consistent with the ground-truth and does not cause confusion.
Please give a score to the following extracted answer.
Appendix D Experiment Details
Our experiments are conducted on three A800 GPUs. All deep models, including the extraction model and decision model, are implemented using the PyTorch framework. For all parts involving the LLMs, we provide 3 examples in the prompt, directly using LLMs to extract without fine-tuning with labeled data. The time out for all LLMs is set to 6s. As for the RL training of the decision module, we set the buffer size to 5000 and the target network update step to 20. For the reward model, we choose Qwen1.5-72B for both Chinese and English datasets.
Hyper-parameters LLMs Decision Model
Batch size - 32
Max length 2048 512
Learning rate - 1e-4
Temperature 0 -
Top p 1 -
Table 3: Hyper-parameters for LLMs and decision model.
To reduce time consumption, for each dataset, we randomly select 2000 items from the training set for the training of the decision model. We train 10 epochs for all datasets. Additionally, the exploration parameter $\epsilon$ is initialized at 0.9 and the discount factor $\gamma$ is set to 0.5. The exploration rate $\epsilon$ becomes 0.9 times its own in every 100 steps until it reaches 0.05.
We employed AdamW optimizer as the optimizer, with using linear scheduling with warming up proportion 10%. For testing, we randomly select 800 items from the test set for each dataset, and the test data is the same for all methods and LLMs. More details are listed in Table 3. Our codes and prompts can be found at https://anonymous.4open.science/r/Ordered-LLM-IE-E89D.
SKE21 DuEE ACE05
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo 88.68 91.12 89.88 87.92 93.57 90.66 91.93 92.50 92.21
Qwen1.5-14B 91.49 93.93 92.69 87.69 88.58 88.13 85.25 89.00 87.08
mistral7B-instruct-v0.3 79.05 88.42 83.47 81.40 87.43 84.31 88.91 90.16 89.53
SKE21 DuEE
Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo + ChatIE 17.37 28.56 21.60 42.31 65.38 51.37
+ ordered -random 19.07 33.87 24.40 55.48 80.57 65.71
+ ordered -fixed 6.62 11.75 8.47 48.16 67.14 56.09
+ ordered -RL (ours) 24.58 40.09 30.47 59.39 86.62 70.47
+24.88% +7.24%
Qwen1.5-14B + ChatIE 25.31 14.06 18.07 48.72 23.75 31.93
+ ordered -random 55.44 63.36 59.14 60.23 67.52 63.66
+ ordered -fixed 49.79 55.76 52.61 57.27 61.46 59.29
+ ordered -RL (ours) 63.69 72.35 67.75 65.19 79.94 71.82
+14.56% +11.36%
Appendix E Additional Results
Here we present some experimental results as a supplement to Section 4. Table 4 shows the results of relation/event classification of different LLMs, and table 5 presents the results of complicated scenarios on SKE21 and DuEE.

### v2

@@ Methods
Furthermore, for different datasets of the same language and task type (such as different English RE datasets), the state structures of them are the same, so their environments are similar. Therefore, only need to mix the training data to train one decision model, which simplifies the training process of the decision model and improves the generalization of our method. The training process is illustrated in Algorithm 1.
Experiments
SKE21 DuEE ACE05
LLM Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT-3.5 88.68 91.12 89.88 87.92 93.57 90.66 91.93 92.50 92.21
Qwen-1.5 91.49 93.93 92.69 87.69 88.58 88.13 85.25 89.00 87.08
Mistral-v0.3 79.05 88.42 83.47 81.40 87.43 84.31 88.91 90.16 89.53
@@
LLM Methods Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT-3.5 ChatIE 81.61 48.51 60.85 71.67 43.31 53.99 70.13 66.13 68.07 57.05 78.13 65.95
LLM + Filter 22.34 25.89 23.98 43.08 51.27 46.82 58.38 72.56 64.70 - - -
@@
Qwen-1.5 ChatIE 55.99 19.85 29.31 48.94 35.25 40.99 47.40 39.14 42.87 49.19 38.06 42.91
LLM + Filter 22.34 25.89 23.98 39.86 57.11 46.95 26.42 48.08 34.10 - - -
@@
Mistral-v0.3 ChatIE 34.56 7.09 11.76 31.50 8.73 13.67 42.84 31.18 36.09 37.73 11.94 18.14
LLM + Filter 6.03 14.37 8.49 28.81 38.60 32.99 36.53 37.27 36.90 - - -
@@
Datasets
We evaluate our methods on several public and accessible complicated information extraction datasets, including the relation extraction datasets NYT (Riedel, Yao, and McCallum 2010), NYT11-HRL (Takanobu et al. 2019), Wiki80 (Han et al. 2019), SKE21 (Xie et al. 2021), HacRED (Cheng et al. 2021) and DuIE (Li et al. 2019), and the event extraction dataset DuEE (Li et al. 2020) and ACE05333https://catalog.ldc.upenn.edu/LDC2006T06. These datasets are challenging for LLMs and widely used for evaluation of existing extraction methods. A brief introduction to these datasets is provided in Technical Appendix.
Comparing Methods and Metrics
We select some recently popular and high-performance LLMs as extractors in our experiments. We choose the open-source model Mistral-7B-instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Qwen1.5-14B (Bai et al. 2023), and the famous GPT3.5-turbo (Brown et al. 2020; Ouyang et al. 2022) for both English and Chinese datasets. We apply two previous multi-stage extraction methods to our testset, ChatIE (Wei et al. 2023) and LLM(without fine-tuning)+Filter (Ding et al. 2024) and compare our method with them. We also compare our method with random/fixed selection of actions, and specifically calculate the metrics in cases involving multiple triples or roles.
We design a relaxed evaluation method based on the exact match. We consider the correctness of the arguments at the token level and calculate the similarity between the extraction result and ground truth for each argument (according to Sec. The Reward Module). In the level of the entire triple or event, we believe the extraction result matches the ground truth only if the relation/event type is classified correctly and the text similarity is greater than the threshold for all arguments. We calculate and report precision (Prec.), recall (Reca.), and F1 scores for all the experiments according to the method above. We repeat the experiment 3 times and averaged the metrics as final results. See Technical Appendix for details.
Effectiveness of LLM Classification
We first assess the efficacy of LLM-based relation/event classification on the SKE21, DuEE and ACE05 datasets. Table 1 shows that LLMs classify the relation/event accurately and achieve stable and accurate results, setting the stage for the argument extraction.
@@
LLM Methods Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT-3.5 ChatIE 46.27 42.70 44.41 32.12 42.97 36.76 32.14 29.57 30.80 57.52 22.18 32.02
LLM + Filter 39.12 44.06 41.44 33.06 47.19 38.88 17.08 18.15 17.60 - - -
@@
Qwen-1.5 ChatIE 39.66 15.66 22.45 27.48 19.46 22.78 16.67 18.29 17.44 41.38 12.29 18.95
LLM + Filter 12.36 19.22 15.04 25.82 29.97 27.74 6.64 9.50 7.82 - - -
@@
Mistral-v0.3 ChatIE 25.61 11.84 16.19 19.38 13.51 15.92 6.18 9.14 7.37 51.16 47.51 49.27
LLM + Filter 38.50 59.91 46.88 16.67 19.41 17.94 8.34 9.92 9.06 - - -
@@
Main Results
Table 2 and  3 show the extraction results on the Chinese and English datasets respectively. ChatIE (Wei et al. 2023) uses a multi-step instruction to guide ChatGPT to extract, we also apply this instruction template to the open source model and report the results, as shown in the first row of each block. We find that the original paper uses zero-shot to conduct experiments on ChatGPT, but for open-source LLMs other than ChatGPT, such zero-shot prompts produced poor results. Therefore, we used the multi-stage prompt template of ChatIE, but added the same number of in-context examples as our method to conduct baseline experiments. Ding et al. (2024) adds the filtered results of small models to the prompt of LLMs [^5] Ding et al. (2024) use the large and small models to collaboratively perform complex triple extraction tasks. They trained a small model that can filter out ”truly related entity pairs” and prompted the filtered results to the large model., and we report the performance of this method on our test set, as shown in the second row of each block. None of the LLMs in the table are fine-tuned on a labeled training set.
@@
SKE21 DuEE
LLM Methods Prec. Reca. F1 Prec. Reca. F1
GPT-3.5 ChatIE 17.37 28.56 21.60 42.31 65.38 51.37
Ordered -random 19.07 33.87 24.40 55.48 80.57 65.71
Ordered -fixed 6.62 11.75 8.47 48.16 67.14 56.09
Ordered -RL (ours) 24.58 40.09 30.47 59.39 86.62 70.47
+24.88% +7.24%
Qwen-1.5 ChatIE 25.31 14.06 18.07 48.72 23.75 31.93
Ordered -random 55.44 63.36 59.14 60.23 67.52 63.66
Ordered -fixed 49.79 55.76 52.61 57.27 61.46 59.29
Ordered -RL (ours) 63.69 72.35 67.75 65.19 79.94 71.82
+14.56% +11.36%
Results of Different Step Orders
To verify the effectiveness of our decision module, we also compared the extraction results of different order selection methods for entity extraction, using Qwen for Chinese data and Mistral for English data, as shown in Table 2 and Table 3. The term ”-random” means randomly selecting actions when extracting arguments, and ”-fixed” means using the same order for all actions (e.g., subject first, then object). In most cases, RL-based extraction order can achieve the best results, and the metrics of fixed extraction order are lower than random order. These results imply that the optimal order varies across distinct instances, and our method could adaptively make the best plan.
Complicated Extraction Settings
@@
The results of Table 4 show that when the IE task becomes complicated, the extraction performance of ChatIE is far inferior to our method. In addition, compared with random or fixed extraction order, our method achieves significant improvements in the metrics. This also suggests that when LLMs undertake extraction tasks in complex scenarios, an appropriate order can further augment the model’s extraction ability.
Conclusion